289500 coulombs were required to deposit one mole of a mettalic element Q from its aqueous solution. calculate the valency of Q
1
Expert's answer
2020-01-14T03:48:25-0500

Faraday's law: m=kq= q*M/(F*z); m =n*M=> z=q/(F*n); n =1mol (condition) F - constant= 96 485.333 C/mol

z=(289500/96485.333*1)=3

valency=3